Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Juri Roda at Sun Feb 11 13:09:27 MET 2001) Success
regular session, no problem

Effelsberg (A. Kraus at Mon Feb 12 09:47:06 MET 2001) Minor failures
Because of a defective receiver control unit, the tsys measurements were corrupted until UT1040. Between UT1000-1048 the antenna was stopped (high wind).

Torun (K.Borkowski at Mon Feb 12 09:51:42 MET 2001) Success
No known problems.

Onsala (John Conway at Mon Feb 12 10:06:20 MET 2001) Minor failures
High Tsys from 0430 due to wet snow in feed. Antenna stowed from 07:15 due to high winds.

Jodrell_Bank (Alastair Gunn at Mon Feb 12 13:20:55 MET 2001) Success
No known problems.

Westerbork (Tony Foley at Mon Feb 12 15:12:56 MET 2001) Success
no known problems

Noto (C.Trigilio at Thu Feb 15 18:36:46 MET 2001) Success
No known problems.